OpenStudy (anonymous):

If f(x) = 2x + 3 and g(x) = x2 + 1, find f(g(3))

OpenStudy (whpalmer4):

Find \(g(3)\). Then use that as \(x\) in \(f(x)\)

OpenStudy (anonymous):

can u pls just type the answer pls its timed and i have 45 sec left pls

OpenStudy (whpalmer4):

Can't help you with a test. Just do the work!

OpenStudy (anonymous):

its not a test its a pre amatory

OpenStudy (whpalmer4):

Whataever it is, the goal is to measure YOUR knowledge and skill, not mine.

OpenStudy (anonymous):

omg cmon pls is it 90?

OpenStudy (anonymous):

there is 23, 47, 82, 90 i think its 47?

OpenStudy (whpalmer4):

What is \((3)^2=\) Now add 1 to that, what do you get? That is \(g(3)\)

OpenStudy (anonymous):

3^2 is 9 and plus one is 10

OpenStudy (whpalmer4):

Good! Now what is \(f((10)=2(10)+3=\)

OpenStudy (anonymous):

ok so the answer is 23 ?

OpenStudy (whpalmer4):

You tell me

OpenStudy (anonymous):

well i think so cuz 20 times 2 is 20 plus 3 is 23

OpenStudy (whpalmer4):

Very good. That is correct. Wasn't so hard, was it?

OpenStudy (anonymous):

not really man thxs :) ARE U A TEACHER?

OpenStudy (whpalmer4):

The key thing to realize here is that x doesn't always mean the same thing. Think of the function as a recipe.

OpenStudy (whpalmer4):

I'm an engineer, but I enjoy showing people what they can do if they are willing to try :-)

OpenStudy (whpalmer4):

So the problem could really be thought of like this: \[a=g(x)\]\[f(a)\]